A DVD (radius 6.0 cm) is spinning freely with an angular velocity of 1150 rpm when a bug drops onto and sticks to the DVD a distance 4.4 cm from the center. If the DVD slows to 900 rpm, what is the ratio of the bugs mass to the DVDs mass? (Ignore the effect of the hole in the center of the DVD.) mbug mOVD
Answer #1

let mass of dvd be mD and that of bug be mB

The problem can be solved using conservation of angular momentum

initial angular velocity = 1150*2*pi/60=120.4 rad/s

final angular velocity = 900*2*pi/60 = 94.25 rad/s

moment of inertia of disc = 0.5*mD*0.062 = 0.0018mD

moment of inertia of bug = mB*0.0442 = 0.0019mB

initial angular momentum =0.0018mD* 120.4 = 0.2167mD

final angular momentum = (0.0018mD+0.0019mB)*94.25= = 0.1696mD+0.1791mB

From conservation of angular momentum,

0.2167mD =  0.1696mD+0.1791mB

0.0471mD = 0.1791mB

mB/mD = 0.0471/0.1791=0.26